# Detecting and Quantifying Isolated Singularities over Discrete Valuation Rings

### v1

@@ 2 Determinacy over Complete Local Rings
As in Section 3.1 of [boubakri2009hypersurface], we can show an analogous results to those of Theorem 2.4 and of Corollary 2.22 for "right equivalence with respect to II". That is, we can say that $f\in R[[\underline{x}]]$ is $k$-right determined with respect to $I$ if for every $g\in R[[\underline{x}]]$ with $f-g\in I^{k+1}$ there exists some $\varphi\in\mathcal{D}_{I}$ such that $\varphi(f)=g$, and we say that $f$ is finite right determined if $f$ is $k$-right determined for some $k$. Therefore, a similar proof to that of Theorem 2.4 tells us that if $I^{k+1}\subset I\cdot J(f)$ then $f$ is $(2k-\operatorname{ord}_{I}(f)+2)$-right determined with respect to $I$, and a similar proof to that of Corollary 2.22 would tell us that if $I\subset\sqrt{J(f)}$ then $f$ is finite right determined.
@@
    If we replace $R$ by a field $k$ and select $I=\langle x_{1},\dots,x_{n}\rangle$, then we have that $j(f)$ would equal to $\tau(f)-1$, where $\tau(f)=\dim_{k}(\frac{k[[\underline{x}]]}{\langle f\rangle+J(f)})$ is the Tjurina number of $f$ over $k$. In fact, we see a similar property in Propostion 3.18. For more information on the Tjurina number of fields and its connection to the topology and algebra of hypersurface singularities, see [greuel2007introduction, milnor2016singular, tjurina1969locally, varchenko1985singularities].
@@
    Note that the definition and the notation of the Jacobian number $j_{I}(f)$ is inspired by the work of Siersma in [siersma1983isolated], Pellikaan in [pellikaan1990deformations], and Svoray in [svoray2024invariants].
@@
Since we are looking at any ideal $I\subset R[[\underline{x}]]$, the enumerative behavior of $j_{I}$ becomes very complicated the more "complicated" the ideal is $I$. For example, as an analogue of Morse’s Lemma (see Theorem 2.46 in [greuel2007introduction]) and of Remark 2.16 in [pellikaan1990deformations]) one expect that (in the case where the characteristic of $\frac{R}{\mathfrak{m}}$ is not $2$) if $I=\langle a_{1},\dots,a_{n}\rangle$ and $f=a_{1}^{2}+\cdots+a_{n}^{2}$ then we have that $J(f)=I$, yet, since $\partial_{i}(f)=\sum_{j=1}^{n}2a_{j}\partial_{i}(a_{j})$, then by Lemma 4.2 in [greuel2019finite], this is true if and only if the Jacobian matrix $D\psi\in\operatorname{Mat}_{r}(V[[\underline{x}]])$ of the $R$-homorphism $\psi$ of $R[[\underline{x}]]$ defined by $\psi(x_{i})=a_{i}$ for every $i$ is invertible, which by Lemma 2.13, this is true if and only if $\psi$ is an automorphism, which only happens when $I=\mathfrak{M}$.
We can use the determinacy theorem with respect to an ideal to conclude a version of the Mather-Yau theorem over $R[[\underline{x}]]$ with respect to an ideal for finitely determined elements. For a zero characteristic version of the Mather-Yau theorem, see Theorem 2.26 in [greuel2007introduction] or [mather1982classification], and for a positive characteristic version, see [greuel2013mather].
@@
Assume that $f\overset{I}{\sim}g$. Then there exists a unit $u\in R[[\underline{x}]]$ and an automorphism $\varphi\in\mathcal{D}_{I}$ of $R[[\underline{x}]]$ such that $g=u\varphi(f)$. Then by Lemma 2.18 we have that $\langle g,I^{k}J(g)\rangle=\langle u\varphi(f),I^{k}J\left(u\varphi(f)\right)\rangle=\langle\varphi(f),I^{k}J\left(\varphi(f)\right)\rangle=\varphi\left(\langle f,I^{k}J(f)\rangle\right)$ for every $k$, which gives us the desired isomorphism of $R$-rings.
Assume that for every big enough $k\gg 0$ we have an isomorphism of $R$-rings $\frac{R[[\underline{x}]]}{\langle f,I^{k}\cdot J(f)\rangle}\cong\frac{R[[\underline{x}]]}{\langle g,I^{k}\cdot J(g)\rangle}$. Therefore, by the Lemma 2.14, this isomorphism lifts to an automorphism $\varphi\colon R[[\underline{x}]]\to R[[\underline{x}]]$ such that $\langle g,I^{k}J(g)\rangle=\varphi\left(\langle f,I^{k}J(f)\rangle\right)$. By Lemma 2.18 we have that $\langle\varphi(f),I^{k}J\left(\varphi(f)\right)\rangle=\varphi\left(\langle f,I^{k}J(f)\rangle\right)$, and so we can assume (up to $\overset{I}{\sim}$) that $\langle g,I^{k}J(g)\rangle=\langle f,I^{k}J(f)\rangle$. Therefore, there exists some $H_{1}\in R[[\underline{x}]]$ and $H_{2}\in I^{k}J(g)$ such that $f=H_{1}g+H_{2}$. By Lemma 2.9 we have that $J(f)\subset I^{\operatorname{ord}_{I}\left(f\right)-1}$ and so $I^{k}J(f)\subset I^{k+\operatorname{ord}_{I}\left(f\right)-1}\subset I^{\operatorname{ord}_{I}\left(f\right)}$. Therefore $g\in I^{\operatorname{ord}_{I}\left(f\right)}$ and $H_{2}\in I^{\operatorname{ord}_{I}\left(f\right)+k-1}$.
@@
Note that the fact that $\langle f\rangle+J(f)=\langle g\rangle+J(g)$ does not give us that $f\overset{\mathcal{K}}{\sim}g$, as in the version of Mather-Yau theorem over $\mathbb{C}$ (see, for example, Theorem 2.26 in [greuel2007introduction]). If the characteristic of $R$ is $p>0$, set $f=x_{1}^{p+1}+x_{2}^{p+1}$ and $g=x_{1}^{p}+x_{1}^{p+1}+x_{2}^{p+1}$. Then we have that $\langle f\rangle+J(f)=\langle g\rangle+J(g)$ but $f\not\overset{\mathcal{K}}{\sim}g$ as they have different orders.
We end this section with a discussion on unfoldings over $R$. Unfoldings play a crucial role in singularity theory over fields, and for a deeper discussion see Chapter II of [greuel2007introduction] or Section 2 of [greuel2016right].
@@
    Note that the results in Proposition 2.31 are much stronger than their counterparts over a field. That is, if $f_{t}$ is a unfolding of $f\in k[[\underline{x}]]$ (for some complete topologized field $k$), then in a neighborhood $U$ of $t=0$ we can only guaranty the inequalities $\mu\left(f_{0}\right)\geq\mu\left(f_{t_{0}}\right)$, $\tau\left(f_{0}\right)\geq\tau\left(f_{t_{0}}\right)$, and $\operatorname{ord}\left(f_{0}\right)\leq\operatorname{ord}\left(f_{t_{0}}\right)$, for every $t_{0}\in U$ (where $\tau$ and $\mu$ are the Tjurina and Milnor number over $k$, respectively). This strengthened results come from the fact that over $R$, the variable $t$ can affect the order of $f_{t}$ but over a field they do not. For more information on the field case, see Appendix A in [greuel2016right].
@@
    Inspired by Section 3 of [greuel2017singularities] and Section I 2.4 of [greuel2007introduction], one can define that $f\in R[[\underline{x}]]$ has finite deformation type if there exists a finite collection of elements $g_{1},...,g_{l}\in R[[x]]$ such that for every unfolding $f_{t}\in R[[x]][t]$ of $f$ there exists some $k$ such that for every $t_{0}\in\mathfrak{m}^{k}$ there exists some $i$ such that $f_{t_{0}}\overset{\mathcal{K}}{\sim}g_{i}\left(x\right)$. In the field case we know that finite deformation type singularities are exactly $ADE$ (for more information on $ADE$ and deformations, see [greuel1990simple, greuel2017singularities]). But over a general local ring, Proposition 2.31 gives us in fact that every element of $R[[\underline{x}]]$ of finite determinacy with respect to $\mathfrak{M}$ has finite deformation type, which by Corollary 2.22, is true for every $f\in R[[\underline{x}]]$ such that $\sqrt{\langle f\rangle+J(f)}=\mathfrak{M}$.
@@ 3 Power Series over a DVR
In this section we focus on on the ring of power series $(V[[\underline{x}]],\mathfrak{m}_{V},\kappa)$ over a DVR $V$ with a chosen uniformizer $\pi$ (where $\mathfrak{m}_{V}=\langle\pi,x_{1},\dots,x_{n}\rangle$). The goal of this section is to develop an analogue of the Tjurina and the Milnor number and of $\overset{\mathcal{K}}{\sim}$ over $V[[\underline{x}]]$ and to show that they have similar properties by "viewing $\pi$ as a variable", inspired by the discussion presented in the Introduction. We formalize this idea using the following Notational setup:
@@
Recall that given an ideal $I$, we say that an ideal $J\subset I$ is a minimal reduction of $I$ if there exists some $N$ such that $JI^{N}=I^{N+1}$ and if there is no ideal $J_{1}\subsetneq J$ with this property (For more information, see Chapter 8 in [huneke2006integral]).
@@
Let $I\subset V[[\underline{x},y]]$ be an ideal such that $\frac{V[[\underline{x},y]]}{I}$ is a one dimensional ring. Then:
@@
    If $\langle a\rangle\,\operatorname{mod}\,I$ is a minimal reduction of the maximal ideal of $\frac{V[[\underline{x},y]]}{I}$ then $\textup{length}_{V[[\underline{x},y]]}\left(\frac{V[[\underline{x},y]]}{I+\langle a\rangle}\right)$ equals to the multiplicity of $\frac{V[[\underline{x},y]]}{I}$ (and therefore is finite and does not depend on the choice of $a$).
@@
The first part is a special case of Theorem 8.6.6 in [huneke2006integral] (which in turn cites [northcott1954reductions, trung2003constructive]). The second part is a special case of Proposition 11.2.2 in [huneke2006integral]. The third part is a special case of semicontinuity of length (for more information, see [504741]). ∎
@@
    Note that if $\mu_{V}\left(f\right)<\infty$ then $\tau_{V}\left(f\right)<\infty$ since $\tau_{V}\left(f\right)\leq\mu_{V}\left(f\right)$ (for a generic choice of $a$). But as in the positive characteristic case (see, for example, Remark 1.2.4. in [boubakri2009hypersurface]), the other direction is note true. Take, for example, $f\left(x\right)=x^{p}+p^{p-1}$ in the case where $p$ is the uniformizer. Yet, as in Proposition 2.1 of [hefez2019hypersurface], we have that if $\tau_{V}\left(f\right)<\infty$ then $\mu_{V}\left(f\right)<\infty$ if and only if $\tilde{f}\in\sqrt{J(\tilde{f})+\langle a\rangle}$ for a generic $a$ as in Definition 3.4.
@@
    We do not set the multiplicity of $\frac{V[[\underline{x}]]}{\langle\tilde{f}\rangle+J(\tilde{f})}$ (resp. of $\frac{V[[\underline{x}]]}{J(\tilde{f})}$) as the definition of $\tau_{V}\left(f\right)$ (resp. $\mu_{V}(f)$) since the multiplicity of $\frac{V[[\underline{x},y]]}{I}$ will always be finite (for every ideal $I$), and as we see in Proposition 3.6, the finiteness of $\tau_{V}\left(f\right)$ as defined above tells us that $f$ has a special property that resembles having an isolated singularity. Note that over a field of positive characteristic $k$, one can define an alternative version of the Milnor number of $f\in k[[\underline{x}]]$ as the multiplicity of $\frac{k[[\underline{x}]]}{J(f)+\langle f\rangle}$, and it behaves much more like its characteristic zero analogue. For more information, see Section 3 of [hefez2019hypersurface].
@@
The following lemma gives us that $\tau_{V}$ detects a property of $f$ that resembles "being an isolated singularity" (but not fully), as a mixed characteristic analogue of Lemma 2.3 in [greuel2007introduction]. We discuss how to detect isolated singularities using an analogue of $\tau_{V}$ in depth in Section 4.
@@
For the second part, if $\tau_{V}\left(f\right)<\infty$ then by Remark 3.5 we have that $\frac{V[[\underline{x},y]]}{\langle\tilde{f}\rangle+J(\tilde{f})}$ must be one dimensional. Thus, since $\frac{V[[\underline{x},y]]}{\tilde{\mathfrak{q}}}$ is two dimensional, we must have that $\langle\tilde{f}\rangle+J(\tilde{f})\not\subset\tilde{\mathfrak{q}}$, and so by the previous part we must have that $\left(\frac{V[[\underline{x},y]]}{\langle f\rangle}\right)_{\tilde{\mathfrak{q}}}$ is regular. ∎
@@
The following proposition allows us to relate between equivalence and "change of variables" (especially in the case where we set $g=\tilde{g_{1}}$ for some $g_{1}\in V[[\underline{x}]]$), as viewed, for example, in the classification process of Theorem B in [carvajal2019covers] and in the notion of "can be written as", as presented in Definition 3.10 of [svoray2025ade].
@@
Given $f\in V[[\underline{x}]]$ and $g\in V[[\underline{x},y]]$, if $\tilde{f}\overset{\mathcal{K}}{\sim}g$ then there exists a sequence $a_{1},\dots,a_{n+1}$ that generates $\mathfrak{m}_{V}$ and some $H\in V[[z_{1},\dots,z_{n+1}]]$ (the ring of formal power series in some dummy variables $z_{1},\dots,z_{n+1}$) whose coefficients are units such that
@@
In addition, if $f\left(\underline{x}\right)=\sum_{\alpha}u_{\alpha}\pi^{n_{\alpha}}\underline{x}^{\alpha}$ as in Notation 3.1 then we can choose $H\left(z_{1},\dots,z_{n+1}\right)=\sum_{\alpha}u_{\alpha}z_{n+1}^{n_{\alpha}}\underline{z}^{\alpha}$ (where $\underline{z}=(z_{1},\dots,z_{n})$).

### v2

@@ 2 Determinacy over Complete Local Rings
Theorem 2.4 generalizes Proposition 2.2 in [CKM22]. In fact, following the notations of their paper, Theorem 2.4 in fact gives us that if $f$ is $k$-determined with respect to $I$ and $\tilde{\alpha}(f)\geq\textup{lct}(\langle f\rangle+I^{k})$.
We now turn to proving a variant of the reverse direction of Corollary 2.22. In order to do so, we recall a version of the local Bertini theorem for complete local rings, as proven by Fenner in [FLE77] for equi-characteristic rings and by Trivedi in [VIJ94] for mixed-characteristic rings. For more information, see Lemma 3.10 in [BAI26] and [OS15].
Proposition 2.24 (Flenner-Trivedi).
Let $A$ be a local Noetherian ring and let $I\subset A$ be an ideal. Then there exists some $x\in I$ such that $x\notin\mathfrak{p}^{(2)}=\mathfrak{p}^{2}R_{\mathfrak{p}}\cap R$ for every $I\not\subset\mathfrak{p}$,
Remark 2.25.
  1. 1.
    Recall that given some $\mathfrak{p}\subset A$ prime ideal and some $f\in A$ (where $A$ local) we have that $\left(\frac{A}{\langle f\rangle}\right)_{\mathfrak{p}}$ is regular if and only if $f\notin\mathfrak{p}^{(2)}$. In addition, over $R[[\underline{x}]]$ we have that if $f\in\mathfrak{p}^{(2)}$ then $J(f)\subset\mathfrak{p}$. For more information on symbolic power of an ideal, see [DDG+15].
  2. 2.
    Proposition 2.24 is not true if we look for an element $x\in I$ such that $x\notin\mathfrak{p}^{(2)}$ for every $I\subset\mathfrak{p}$. For example, if $I=\mathfrak{q}^{2}$ for some prime ideal $\mathfrak{q}$ then $I$ must be contained in $\mathfrak{p}^{(2)}$ for every prime ideal $I\subset\mathfrak{p}$, and so given some $x\in I$, then $x\in\mathfrak{p}^{(2)}$ for every prime ideal $I\subset\mathfrak{p}$.
  3. 3.
    Recall that the singular locus of $\frac{R[[\underline{x}]]}{\langle f\rangle}$, denoted $\textup{Sing}(V(f))$, is the set of prime ideals $\mathfrak{p}\subset R[[\underline{x}]]$ for which $\left(\frac{R[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{p}}$ is a regular local ring. By the previous item, $\textup{Sing}(V(f))$ is exactly the set of all prime ideals for which $f\in\mathfrak{p}^{(2)}$ and $\textup{Sing}(V(f))\subset V(J(f)+\langle f\rangle)$. In addition, from Lemma 2.18 we can conclude that if $\varphi(f)=g$ for some $R$-automorphism of $R[[\underline{x}]]$ then $\varphi^{*}(\textup{Sing}(V(f))=\textup{Sing}(V(g))$. For more information on the singular locus of a ring, see Sections 15.47 and 15.48 in [STA18].
Proposition 2.26.
Let $I\subset R[[\underline{x}]]$ be an ideal and let $f\in R[[\underline{x}]]$ be finite determined with respect to $I$. Then $\textup{Sing}(V(f))\subset V(I+\langle f\rangle+J(f))$.
Proof.
Assume that $f\in R[[\underline{x}]]$ is $k$-determined with respect to $I$ (up to increasing $k$ we can assume that $k>0$). By Proposition 2.24 there exists some $g\in I^{k+1}$ such that $f+g\notin\mathfrak{p}^{(2)}$ for every prime ideal $\langle f\rangle+I^{k+1}\not\subset\mathfrak{p}$.
We show that $\textup{Sing}(V(f+g))\subset V(I+J(f)+\langle f\rangle)$. Given some prime ideal $\mathfrak{p}$ such that $I+J(f)+\langle f\rangle\not\subset\mathfrak{p}$, we have two options: First, if $I\not\subset\mathfrak{p}$ then $\langle f\rangle+I\not\subset\mathfrak{p}$ and so by the choice of $g\in I$ we must have that $f+g\notin\mathfrak{p}^{(2)}$. Second, if $I\subset\mathfrak{p}$ then as $I+J(f)+\langle f\rangle\not\subset\mathfrak{p}$ we must have that $J(f)+\langle f\rangle\not\subset\mathfrak{p}$ and so by the first item of Remark 2.25 we have that $f\notin\mathfrak{p}^{(2)}$ and so $f+g\notin\mathfrak{p}^{(2)}$. Therefore, from the third item of Remark 2.25, we can conclude that if $\mathfrak{p}\in\textup{Sing}(V(f+g))$ then $I+J(f)+\langle f\rangle\subset\mathfrak{p}$.
Since $f\in R[[\underline{x}]]$ is $k$-determined with respect to $I$ then $f+g\overset{I}{\sim}f$ and so there exists some $\varphi\in\mathcal{D}_{I}$ such that $\langle\varphi(f)\rangle=\langle f+g\rangle$. By Lemma 2.18 we have that $\varphi(I+J(f)+\langle f\rangle)=I+J(f+g)+\langle f+g\rangle$ (as $\varphi\in\mathcal{D}_{I}$). Since $g\in I^{k+1}$ then from Lemma 2.9 we have that $\partial_{i}(g)\in I^{k}\subset I$ and therefore $I+J(f)+\langle f\rangle=I+J(f+g)+\langle f+g\rangle$. Thus, by the third item of Remark 2.25, we can conclude that $\textup{Sing}(V(f))\subset V(I+J(f)+\langle f\rangle)$ as desired. ∎
Theorem 2.27.
Assume that $\textup{Sing}(V(f))=V(J(f)+\langle f\rangle)$. Then $f$ has finite determinacy with respect to $I$ if and only if $I\subset\sqrt{\langle f\rangle+J(f)}$.
Proof.
This follows directly from Corollary 2.22 and Proposition 2.26, recalling that $V(I+J(f)+\langle f\rangle)=V(I)\cap V(J(f)+\langle f\rangle)$. ∎
Remark 2.28.
  1. 1.
    The assumption $\textup{Sing}(V(f))=V(J(f)+\langle f\rangle)$ in Theorem 2.27 is true for every $f\in R[[\underline{x}]]$ when $R$ is equi-characteristic (i.e. contains a field) and this follows from the classical Jacobian criterion (see, for example, Theorem 6.28 in [GW10]). In particular, in the equicharacteristic case, we can view Theorem 2.27 as a generalization of Theorem 2.8 in [BGM12] to hypersurfaces with arbitrary singularities (noting that their proof relies on a careful analysis of the action of the automorphism group). Yet, as we see in Theorem 4.8 and in Proposition 4.24, we might have that $\textup{Sing}(V(f))\subsetneq V(J(f)+\langle f\rangle)$ if $R$ is mixed characteristic.
  2. 2.
@@
    As in Section 3.1 of [BOU09], we can show an analogous results to those of Theorem 2.4 and of Corollary 2.22 for "right equivalence with respect to II". That is, we can say that $f\in R[[\underline{x}]]$ is $k$-right determined with respect to $I$ if for every $g\in R[[\underline{x}]]$ with $f-g\in I^{k+1}$ there exists some $\varphi\in\mathcal{D}_{I}$ such that $\varphi(f)=g$, and we say that $f$ is finite right determined if $f$ is $k$-right determined for some $k$. Therefore, a similar proof to that of Theorem 2.4 tells us that if $I^{k+1}\subset I\cdot J(f)$ then $f$ is $(2k-\operatorname{ord}_{I}(f)+2)$-right determined with respect to $I$, and a similar proof to that of Corollary 2.22 would tell us that if $I\subset\sqrt{J(f)}$ then $f$ is finite right determined.
@@
    If we replace $R$ by a field $k$ and select $I=\langle x_{1},\dots,x_{n}\rangle$, then we have that $j(f)$ would equal to $\tau(f)-1$, where $\tau(f)=\dim_{k}(\frac{k[[\underline{x}]]}{\langle f\rangle+J(f)})$ is the Tjurina number of $f$ over $k$. In fact, we see a similar property in Propostion 3.18. For more information on the Tjurina number of fields and its connection to the topology and algebra of hypersurface singularities, see [GLS07, MIL16, TJU69, VG85].
@@
    Note that the definition and the notation of the Jacobian number $j_{I}(f)$ is inspired by the work of Siersma in [SIE83], Pellikaan in [PEL90], and Svoray in [SVO24].
@@
  1. 1.
    Item 2 of Proposition 2.31 gives us a quantitative bound on the determinacy of $f$ with respect to $I$. It is analogue of the role of the Milnor and Tjurina numbers in Corollary 2.4 of [BGM12] and of the role of the $c_{I}(f)$ invariant in Theorem 1.6 in [DE 88], Theorem 6.5. in [PEL88], and Proposition 1.6 in [SIE83]).
  2. 2.
    Since we are looking at any ideal $I\subset R[[\underline{x}]]$, the enumerative behavior of $j_{I}$ becomes very complicated the more "complicated" the ideal is $I$. For example, as an analogue of Morse’s Lemma (see Theorem 2.46 in [GLS07]) and of Remark 2.16 in [PEL90]) one expect that (in the case where the characteristic of $\frac{R}{\mathfrak{m}}$ is not $2$) if $I=\langle a_{1},\dots,a_{n}\rangle$ and $f=a_{1}^{2}+\cdots+a_{n}^{2}$ then we have that $J(f)=I$, yet, since $\partial_{i}(f)=\sum_{j=1}^{n}2a_{j}\partial_{i}(a_{j})$, then by Lemma 4.2 in [GP19], this is true if and only if the Jacobian matrix $D\psi\in\operatorname{Mat}_{r}(R[[\underline{x}]])$ of the $R$-homorphism $\psi$ of $R[[\underline{x}]]$ defined by $\psi(x_{i})=a_{i}$ for every $i$ is invertible, which by Lemma 2.13, this is true if and only if $\psi$ is an automorphism, which only happens when $I=\mathfrak{M}$.
We can use the determinacy theorem with respect to an ideal to conclude a version of the Mather-Yau theorem over $R[[\underline{x}]]$ with respect to an ideal for finitely determined elements. For a zero characteristic version of the Mather-Yau theorem, see Theorem 2.26 in [GLS07] or [MY82], and for a positive characteristic version, see [GP13].
@@
Assume that $f\overset{I}{\sim}g$. Then there exists a unit $u\in R[[\underline{x}]]$ and an automorphism $\varphi\in\mathcal{D}_{I}$ of $R[[\underline{x}]]$ such that $g=u\varphi(f)$. Then by Lemma 2.18 we have that
$\langle g,I^{k}J(g)\rangle=\langle u\varphi(f),I^{k}J\left(u\varphi(f)\right)\rangle=\langle\varphi(f),I^{k}J\left(\varphi(f)\right)\rangle=\varphi\left(\langle f,I^{k}J(f)\rangle\right)$
for every $k$, which gives us the desired isomorphism of $R$-rings.
Assume that for every big enough $k\gg 0$ we have an isomorphism of $R$-rings $\frac{R[[\underline{x}]]}{\langle f,I^{k}\cdot J(f)\rangle}\cong\frac{R[[\underline{x}]]}{\langle g,I^{k}\cdot J(g)\rangle}$. Therefore, by the Lemma 2.14, this isomorphism lifts to an automorphism $\varphi\colon R[[\underline{x}]]\to R[[\underline{x}]]$ such that $\langle g,I^{k}J(g)\rangle=\varphi\left(\langle f,I^{k}J(f)\rangle\right)$. By Lemma 2.18 we have that
$\langle\varphi(f),I^{k}J\left(\varphi(f)\right)\rangle=\varphi\left(\langle f,I^{k}J(f)\rangle\right),$
and so we can assume (up to $\overset{I}{\sim}$) that $\langle g,I^{k}J(g)\rangle=\langle f,I^{k}J(f)\rangle$. Therefore, there exists some $H_{1}\in R[[\underline{x}]]$ and $H_{2}\in I^{k}J(g)$ such that $f=H_{1}g+H_{2}$. By Lemma 2.9 we have that $J(f)\subset I^{\operatorname{ord}_{I}\left(f\right)-1}$ and so $I^{k}J(f)\subset I^{k+\operatorname{ord}_{I}\left(f\right)-1}\subset I^{\operatorname{ord}_{I}\left(f\right)}$. Therefore $g\in I^{\operatorname{ord}_{I}\left(f\right)}$ and $H_{2}\in I^{\operatorname{ord}_{I}\left(f\right)+k-1}$.
@@
Note that the fact that $\langle f\rangle+J(f)=\langle g\rangle+J(g)$ does not give us that $f\overset{\mathcal{K}}{\sim}g$, as in the version of Mather-Yau theorem over $\mathbb{C}$ (see, for example, Theorem 2.26 in [GLS07]). If the characteristic of $R$ is $p>0$, set $f=x_{1}^{p+1}+x_{2}^{p+1}$ and $g=x_{1}^{p}+x_{1}^{p+1}+x_{2}^{p+1}$. Then we have that $\langle f\rangle+J(f)=\langle g\rangle+J(g)$ but $f\not\overset{\mathcal{K}}{\sim}g$ as they have different orders.
We end this section with a discussion on unfoldings over $R$. Unfoldings play a crucial role in singularity theory over fields, and for a deeper discussion see Chapter II of [GLS07] or Section 2 of [GN16].
@@
    The results in Proposition 2.36 are much stronger than their counterparts over a field. That is, if $f_{t}$ is a unfolding of $f\in k[[\underline{x}]]$ (for some complete topologized field $k$), then in a neighborhood $U$ of $t=0$ we can only guaranty the inequalities $\mu\left(f_{0}\right)\geq\mu\left(f_{t_{0}}\right)$, $\tau\left(f_{0}\right)\geq\tau\left(f_{t_{0}}\right)$, and $\operatorname{ord}\left(f_{0}\right)\leq\operatorname{ord}\left(f_{t_{0}}\right)$, for every $t_{0}\in U$ (where $\tau$ and $\mu$ are the Tjurina and Milnor number over $k$, respectively). This strengthened results come from the fact that over $R$, the variable $t$ can affect the order of $f_{t}$ but over a field they do not. For more information on the field case, see Appendix A in [GN16].
@@
    Inspired by Section 3 of [GRE17] and Section I 2.4 of [GLS07], one can define that $f\in R[[\underline{x}]]$ has finite deformation type if there exists a finite collection of elements $g_{1},...,g_{l}\in R[[x]]$ such that for every unfolding $f_{t}\in R[[x]][t]$ of $f$ there exists some $k$ such that for every $t_{0}\in\mathfrak{m}^{k}$ there exists some $i$ such that $f_{t_{0}}\overset{\mathcal{K}}{\sim}g_{i}\left(x\right)$. In the field case we know that finite deformation type singularities are exactly $ADE$ (for more information on $ADE$ and deformations, see [GK90, GRE17]). But over a general local ring, Proposition 2.36 gives us in fact that every element of $R[[\underline{x}]]$ of finite determinacy with respect to $\mathfrak{M}$ has finite deformation type, which by Corollary 2.22, is true for every $f\in R[[\underline{x}]]$ such that $\sqrt{\langle f\rangle+J(f)}=\mathfrak{M}$.
@@ 3 Power Series over a DVR
In this section we focus on on the ring of power series $(V[[\underline{x}]],\mathfrak{m}_{V},\kappa)$ over a DVR $V$ with a chosen uniformizer $\pi$ (where $\mathfrak{m}_{V}=\langle\pi,x_{1},\dots,x_{n}\rangle$). The goal of this section is to develop an analogue of the Tjurina and the Milnor number and of $\overset{\mathcal{K}}{\sim}$ over $V[[\underline{x}]]$ and to show that they have similar properties by "viewing $\pi$ as a variable", inspired by the discussion presented in Section 1. We formalize this idea using the following Notational setup:
@@
Recall that given an ideal $I$, we say that an ideal $J\subset I$ is a minimal reduction of $I$ if there exists some $N$ such that $JI^{N}=I^{N+1}$ and if there is no ideal $J_{1}\subsetneq J$ with this property (For more information, see Chapter 8 in [HS06]).
@@
Let $I\subset V[[\underline{x},y]]$ be an ideal such that $\frac{V[[\underline{x},y]]}{I}$ is a one-dimensional ring. Then:
@@
    If $\langle a\rangle\,\operatorname{mod}\,I$ is a minimal reduction of the maximal ideal of $\frac{V[[\underline{x},y]]}{I}$ then we have that $\textup{length}_{V[[\underline{x},y]]}\left(\frac{V[[\underline{x},y]]}{I+\langle a\rangle}\right)$ equals to the multiplicity of $\frac{V[[\underline{x},y]]}{I}$ (and therefore is finite and does not depend on the choice of $a$).
@@
The first part is a special case of Theorem 8.6.6 in [HS06] (which in turn cites [NR54, TRU03]). The second part is a special case of Proposition 11.2.2 in [HS06]. The third part is a special case of semicontinuity of length (for more information, see [EPS25]). ∎
@@
    Note that if $\mu_{V}\left(f\right)<\infty$ then $\tau_{V}\left(f\right)<\infty$ since $\tau_{V}\left(f\right)\leq\mu_{V}\left(f\right)$ (for a generic choice of $a$). But as in the positive characteristic case (see, for example, Remark 1.2.4. in [BOU09]), the other direction is note true. Take, for example, $f\left(x\right)=x^{p}+p^{p-1}$ in the case where $p$ is the uniformizer. Yet, as in Proposition 2.1 of [HRS19], we have that if $\tau_{V}\left(f\right)<\infty$ then $\mu_{V}\left(f\right)<\infty$ if and only if $\tilde{f}\in\sqrt{J(\tilde{f})+\langle a\rangle}$ for a generic $a$ as in Definition 3.4.
@@
    We do not set the multiplicity of $\frac{V[[\underline{x}]]}{\langle\tilde{f}\rangle+J(\tilde{f})}$ (resp. of $\frac{V[[\underline{x}]]}{J(\tilde{f})}$) as the definition of $\tau_{V}\left(f\right)$ (resp. $\mu_{V}(f)$) since the multiplicity of $\frac{V[[\underline{x},y]]}{I}$ will always be finite (for every ideal $I$), and as we see in Proposition 3.6, the finiteness of $\tau_{V}\left(f\right)$ as defined above tells us that $f$ has a special property that resembles having an isolated singularity. Note that over a field of positive characteristic $k$, one can define an alternative version of the Milnor number of $f\in k[[\underline{x}]]$ as the multiplicity of $\frac{k[[\underline{x}]]}{J(f)+\langle f\rangle}$, and it behaves much more like its characteristic zero analogue. For more information, see Section 3 of [HRS19].
@@
The following lemma gives us that $\tau_{V}$ detects a property of $f$ that resembles "being an isolated singularity" (but not fully), as a mixed characteristic analogue of Lemma 2.3 in [GLS07]. We discuss how to detect isolated singularities using an analogue of $\tau_{V}$ in depth in Section 4.
@@
For the second part, if $\tau_{V}\left(f\right)<\infty$ then by Remark 3.5 we have that $\frac{V[[\underline{x},y]]}{\langle\tilde{f}\rangle+J(\tilde{f})}$ must be one-dimensional. Thus, since $\frac{V[[\underline{x},y]]}{\tilde{\mathfrak{q}}}$ is two dimensional, we must have that $\langle\tilde{f}\rangle+J(\tilde{f})\not\subset\tilde{\mathfrak{q}}$, and so by the previous part we must have that $\left(\frac{V[[\underline{x},y]]}{\langle f\rangle}\right)_{\tilde{\mathfrak{q}}}$ is regular. ∎
@@
The following proposition allows us to relate between equivalence and "change of variables" (especially in the case where we set $g=\tilde{g_{1}}$ for some $g_{1}\in V[[\underline{x}]]$), as viewed, for example, in the classification process of Theorem B in [CMP+21] and in the notion of "can be written as", as presented in Definition 3.10 of [SVO25].
@@
Given $f\in V[[\underline{x}]]$ and $g\in V[[\underline{x},y]]$, if $\tilde{f}\overset{\mathcal{K}}{\sim}g$ then there exists a sequence $a_{1},\dots,a_{n+1}$ that generates $\mathfrak{m}_{V}$ and there exists some $H\in V[[z_{1},\dots,z_{n+1}]]$ (the ring of formal power series in some dummy variables $z_{1},\dots,z_{n+1}$) whose coefficients are units such that
@@
In addition, if $f\left(\underline{x}\right)=\sum_{\alpha}u_{\alpha}\pi^{n_{\alpha}}\underline{x}^{\alpha}$ as in Notation 3.1 then for $\underline{z}=(z_{1},\dots,z_{n})$ we can choose $H\left(z_{1},\dots,z_{n+1}\right)=\sum_{\alpha}u_{\alpha}z_{n+1}^{n_{\alpha}}\underline{z}^{\alpha}$.